# BESSTIE: A Benchmark for Sentiment and Sarcasm Classification for Varieties of English

### v1

@@ 7.3. Cross-variety/domain evaluation
Similarly, to assess the cross-domain robustness of our sentiment classification models, we conduct experiments with domain data from google, and reddit. Models trained on one domain are tested on the other domain, and vice-versa, for each of the three varieties of English. Figure 7 shows that models perform better when trained on in-domain data, for both mistral and bert. We observe that pre-trained mistral performs better for all varieties in comparison to results from the cross-domain experiments.
@@
While mistral outperforms bert at both in- and cross-domain experiments, there is significant variation in cross-domain results, highlighting the challenge in generalization, given domain variance in fine-tuned data. Moreover, the direction of cross-domain experiments also affects performance. Both models perform better for reddit $\rightarrow$ google, suggesting that models may be better at transferring sentiment from a relatively informal domain. The overall difference in cross-domain performance is significantly higher for inner-circle varieties, compared to en-IN, indicating that fine-tuning with mixed domain data may be less effective for such varieties. Across varieties, however, en-UK, an inner-circle variety, performs better for google $\rightarrow$ reddit, whereas for the bert model, this is observed with the en-IN variety.
8. Error Analysis
We randomly sample 30 examples misclassified by mistral, across all domains and varieties. We hope that these error categories will be helpful for future researchers who use the BESSTIE dataset. For sentiment classification, we identify the following error categories:
  1. (1)
    Mixed Sentiment Mis-classification: The model struggles to classify text containing both positive and negative sentiment, as in ‘The outside pool has been completely redesigned and although looks good the depth of the pool has been downsized.’.
  2. (2)
    Insufficient Context: Consider the Reddit comment ‘Darwin for getting its own dot and big font on any map of the World.’. The sentiment can only be identified with the context/information from the previous exchanges in the thread.
For sarcasm classification, the error categories are:
  1. (1)
    Contextual Sarcasm: Some instances of sarcasm rely heavily on contextual or conversational cues that are not present in the text on its own. For example, in the Reddit comment, ‘Atleast Spiderman covered his face before answering.’, the sarcastic intent can only be understood within the context of the preceding discussion in the thread.
  2. (2)
    Implicit Sarcasm: Instances where sarcasm is expressed subtly, through tone or mild exaggeration, without explicit contradictions. The example, ‘Makes more sense to just bar his company from getting further contracts whilst he is running it.’, is misidentified as not sarcastic.
  3. (3)
    False Positives from Negative Connotations: Mildly negative statements are often misclassified as sarcastic due to their phrasing or sentiment. This error is identified for examples such as ‘Sorry gal, what you experienced pretty much IS the dating culture here.’.
9. Conclusion
Biases in LLMs have been reported for several language varieties. This paper presents BESSTIE, a benchmark for sentiment and sarcasm classification across three varieties of English: Australian (en-AU), Indian (en-IN), and British (en-UK). Our evaluation spans nine LLMs, six encoders and three decoders, assessed on datasets collected from Google Places reviews (google) and Reddit comments (reddit). The models perform consistently better on en-AU and en-UK (i.e., inner-circle varieties) than on en-IN (i.e., the outer-circle variety) for both tasks. Although models report high performance on sentiment classification (F-Score of 0.81 and 0.75, for google and reddit respectively), they struggle with detecting sarcasm (F-Score of 0.59 on reddit), indicating that sarcasm classification is still largely unsolved. This is supported by our cross-variety evaluation which reveals limited generalisation capability for sarcasm classification, where cultural and contextual understanding is crucial. Notably, monolingual models marginally outperform multilingual models, suggesting that language diversity in pre-training does not extend to varieties of a language. In conclusion, BESSTIE will provide a way to measure bias of LLMs towards non-standard English varieties for the two tasks. By highlighting gaps in the model performance, especially for (a) sarcasm classification in general; and (b) outer-circle varieties like en-IN for both the tasks, this benchmark will be useful to evaluate dialectal biases of language models, and contribute to research towards equitable performance across all varieties of English.
Acknowledgments
This work was funded by Google’s Research Scholar grant awarded in 2024.

### v3

@@ Appendix D Cross-domain Evaluation
We conduct experiments with domain data from google, and reddit, to assess the cross-domain robustness of our sentiment classification models. Models trained on one domain are tested on the other domain, and vice-versa, for each of the three varieties of English. Figure 7 shows that models perform better when trained on in-domain data, for both mistral and bert. We observe that pre-trained mistral performs better for all varieties in comparison to results from the cross-domain experiments.
Figure 6: Cross-variety performance analysis of bert. The figure compares two different scenarios: in-variety fine-tuning, and cross-variety fine-tuning for sentiment and sarcasm classification across all varieties.
The corresponding results for bert are in Figure 6.
Figure 7: Cross-domain performance analysis of mistral and bert comparing a decoder and an encoder model.
@@
While mistral outperforms bert at both in- and cross-domain experiments, there is significant variation in cross-domain results, highlighting the challenge in generalization, given domain variance in fine-tuned data. Moreover, the direction of cross-domain experiments also affects performance. Both models perform better for reddit $\rightarrow$ google, suggesting that models may be better at transferring sentiment from a relatively informal domain. The overall difference in cross-domain performance is significantly higher for inner-circle varieties, compared to en-IN, indicating that fine-tuning with mixed-domain data may be less effective for such varieties. Across the varieties, however, en-UK, an inner-circle variety, performs better for google $\rightarrow$ reddit, whereas for the bert model, this is observed with the en-IN variety.
Appendix E Detailed Feature Types
We present detailed feature types identified in misclassified samples from different English varieties using the mistral model. Each table lists the feature, an example illustrating the feature, and the frequency count. Table 11 shows features found in en-AU samples, Table 12 shows features found in en-IN samples, and Table 13 shows features found in en-UK samples.
Feature Example Count dial levelling of past tense verb forms when we order passionfruit cheecake and never get it 2 pronoun drop (We) also ordered 2 noodles, fish and dessert 3 there’s with plural subjects There’s been plenty of fossils and opals uncovered 1 no auxiliary in yes/no questions (Do) people still watch commercial TV? 2 coll Hi Hospo workers afternoon when you wake up 28 cont Beige night 6
Table 11: Features identified in the en-AU samples misclassified by mistral.
Feature Example Count dial article omission Kamal is playing (a) cameo role 35 pronoun drop (I) can’t wait for animal park man 22 object/subject fronting Laws of India doesn’t apply on white man, he is above everything 12 ‘very’ as qualifier quantity is very very less 10 copula omission (I am) Waiting for cash to arrive in a tempo 18 coll It means benstokes 33 cont Alia Advani as her earlier name suits better. 3 code Bridge chori ho jaaega 8
Table 12: Features identified in the en-IN samples misclassified by mistral.
Feature Example Count dial was for conditional were if it was in the autumn. 1 group genitive their and their old boy network 1 levelling of past tense verb forms but could been more spicy. 1 pronoun drop (It) took just 2 and a half minutes to lose his head 4 coll Warra calm and coherent interview. 15 cont Haha yup! 5
Table 13: Features identified in the en-UK samples misclassified by mistral.